Shifter leader was AG Shifters were 1. Around 6 hrs of beam<br> <br> 2. Fill pattern: 106 X 106_p2, polarization: B/Y: -44.99 +/- 2.33, -47.06 +/- 2.28:<br> Run 15637,15638, 15640<br> The IoServer crashed at 01.55 am in run #15638.<br> <br> New polarization (01.40 am): B/Y: -40.28 +/- 2.38, -47.90 +/- 2.33:<br> Run 15643, 15644<br> <br> New polarization (02.25 am): B/Y: -43.10 +/- 2.39, -52.50 +/- 2.36:<br> Run 15645-15650<br> <br> Fill pattern: 111 X 111_p1, new polarization (07.10 am): B/Y: -48.63 +/- 2.37, -47.24 +/- 2.25:<br> Run 15653,15655<br> <br> Trig2(1213k), Trig3(205k), Trig4(155k) ,Trig5(1141k)<br> <br> Pedestal run 15651, Non pedestal run 15652<br> <br> 3. New error messages (tripped!) first seen around 03.00 am:<br> TPM2: Drift (blue): demand: 30, actual 26<br> T1: Field (blue): demand -4900, actual: -6<br> _______________________________________________ Brahms-l mailing list Brahms-l@lists.bnl.gov http://lists.bnl.gov/mailman/listinfo/brahms-lReceived on Sun Jun 19 08:12:38 2005
This archive was generated by hypermail 2.1.8 : Sun Jun 19 2005 - 08:12:50 EDT